OpenStudy (anonymous):

8.2/0.5 please help im in 5th grade doing 6th grade math work need a little help

OpenStudy (charlotte123):

@supercam what can I help u with? ^_^

OpenStudy (anonymous):

diving by decimals im in 5th grade

OpenStudy (anonymous):

8.2/0.5

OpenStudy (anonymous):

its so hard

OpenStudy (charlotte123):

16.4 :D

OpenStudy (charlotte123):

that is the answer ^_^ http://www.wolframalpha.com/input/?i=8.2%2F0.5

OpenStudy (anonymous):

thank u...how did u get that I got 16.0

OpenStudy (anonymous):

\[\frac{8.2}{0.5}=\frac{82}{100}/\frac{5}{100}=\frac{82}{5}=16.4 \]

OpenStudy (anonymous):

our teacher never taught us that way

OpenStudy (anonymous):

In a fraction, with decimal numbers, you can always shift the decimal point the same number of places in both the numerator and denominator the same number of places either right or left.

OpenStudy (anonymous):

o ok i keep forgetting to do that